Given p(x)=x^3+x^2+x+1. Write P(x) in factored form. Be sure to write the full equation including P(x)?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

-1/1---1---1---1 synthetic division
1 0 1 0
(x+1)(x^2+1)=P(x)

Answer by ikleyn(52802)

.
Apply grouping as the factoring method:


x^3 + x^2 + x + 1 = (x^3 + x^2) + (x + 1) = x^2*(x+1) + (x+1) = (x+1)*(x^2+1).

Solved.

That's all.


-----------------

Regarding your writing, please learn, where and when to use the terms "equation" and "expression".

These terms are different, and correct using the terms is the first thing on how the specialists recognize a newcomer.